OpenStudy (anonymous):

find the value of c such that one solution to 4x^2-6x+c=0 is x=3.

OpenStudy (anonymous):

\[4x^{2}-6x+c=0\]\[4(3)^{2}-6(3)+c=0\]\[4(9)-6(3)+c=0\]\[36-18+c=0\]You finish it from there.

OpenStudy (anonymous):

36-18+c=0 -18+c=0 +18 +18 c=18 from here idk what else to do

OpenStudy (anonymous):

\[36-18+c=0\]\[18+c=0\]\[c=-18\]

OpenStudy (anonymous):

oh yeah idk what i was thinking

OpenStudy (anonymous):

is that it ?

OpenStudy (anonymous):

Yes